In a single-sample t-test, a random sample of n = 25 individuals is selected from a population with μ = 20, and a treatment is administered to each individual in the sample. After treatment, the sample mean is found to be M = 22.2 with SS = 384. Compute the value of the single-sample t-test.
Answer #1

SS the sum of squares = 384. Therefore standard deviation s = Sqrt(SS/n-1) = Sqrt(384/24) = Sqrt(16) = 4

The Test statistic is given by

t = \frac{\bar{x}-\mu }{\frac{s}{\sqrt{n}}} = \frac{22.2-20}{\frac{4}{\sqrt{25}}} = 2.75

The value = 2.75